OpenStudy (anonymous):

Find the exact length of the curve. Use a graph to determine the parameter interval. r=cos^2(theta/2)

OpenStudy (anonymous):

The formula for the length of a polar curve is \[\int\limits_{a}^{b} \sqrt{r^2+(dr/d\theta)^2}\]

OpenStudy (anonymous):

Do you see how to write the integrand?

OpenStudy (anonymous):

Heres what I got using that, which gave me a wrong answer... \[\int\limits_{0}^{4\pi}\sqrt{(\cos ^{4}(\theta/2)+(\sin^{2}(\theta))/4}\]

OpenStudy (anonymous):

I think the second term should just be sin^2(theta)

OpenStudy (anonymous):

\[r=\cos^{2}(\theta/2), r^{2}=\cos^{4}(\theta/2)\]

OpenStudy (anonymous):

No wait, you were right about that derivative

OpenStudy (anonymous):

okay, should i use half angle formula for the r^2 piece under the sqrt?

OpenStudy (anonymous):

You shouldn't need to, you're just evaluating this numerically, right?

OpenStudy (anonymous):

yea, maybe my bounds are incorrect, i used wolfram to get the bounds

OpenStudy (anonymous):

Wolfram tends to be correct, and 4pi does make sense.

OpenStudy (anonymous):

so why is my answer of 8 incorrect?

OpenStudy (anonymous):

Actually, wolfram says it should be 0 to 2pi for me

OpenStudy (anonymous):

yep that was the problem, thanks for the help

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

How did you guys do the integral?